# Security Approaches for Data Provenance in the Internet of Things: A Systematic Literature Review

### v1

@@ 3. Data provenance in IoT networks
3.6   Security Attacks — IoT networks are vulnerable to active and passive attacks. The proposed security techniques takes into account these attacks and try to prove their robustness against it through many proposed solutions using different technologies. Provenance information may be more sensitive than the data itself, which makes it a priority for attackers. Attacks on such systems are of two types: attacks on data and attacks on provenance. Next, we outline some representative attacks that may be perpetrated in an IoT environment against data and provenance.
3.6.1   Data Attacks
@@
  • Packet drop attack: Is a security threat where an adversary intentionally discards or drops specific data packets within the network. This malicious act disrupts communication between IoT devices and can lead to various issues, such as data loss, delayed information delivery, or service unavailability. As IoT devices often rely on wireless communication and may operate in resource-constrained environments, packet drop attacks can significantly impact the overall system’s functionality and reliability.
@@
  • Packet replay attack: Is a type of security attack where an adversary intercepts and records data packets as they travel through the network. The attacker then replays or re-sends these recorded packets at a later time, attempting to deceive the system into thinking the information is fresh and legitimate. This malicious action can lead to various security issues, such as unauthorized access, data manipulation, or false triggering of actions based on the replayed data.
@@
  • Data forgery: Is a form of security attack where an adversary manipulates or alters the data transmitted between IoT devices. The attacker may forge false sensor readings, control commands, or other critical information to deceive the system or cause malicious actions. This type of attack can lead to incorrect decisions, unauthorized access, or compromised system integrity.
@@
  • Data modification attack: Is a type of security attack where an unauthorized entity alters or modifies the data being transmitted between IoT devices. The attacker may tamper with sensor readings, control commands, or other critical information, leading to incorrect decisions or actions within the system. This can result in operational disruptions, potential risks, or unauthorized access to sensitive data.
@@
  • Eavesdrop: Refers to the unauthorized interception and monitoring of communication between IoT devices. A malicious attacker, also known as an eavesdropper, captures and listens to data packets as they traverse the network. By doing so, the eavesdropper can access sensitive information, such as sensor readings, control commands, or personal data, which can lead to privacy breaches, security vulnerabilities, and potential misuse of the intercepted data.
3.6.2   Provenance Attacks
@@ 4. Existing surveys
Several review articles address the issue of data provenance in IoT networks. Most of these articles focus on the security requirements for data provenance in IoT. In this section, we compare and highlight the importance and need to conduct this review paper. Table 3 provides a comparison between the different review articles based on a number of aspects, including time range (represent the starting and ending years of the conduct review), review methodology (Does the article follows a specified review methodology to conduct the review?), taxonomy (Does the authors provide researchers with a taxonomy that summaries the different aspects used in the review?), security requirements (Does the review use and elaborate the needed security requirements in data provenance approaches?), attacks (Does the review analyze the different approaches based on the security attacks?), performance metrics (Does the review analyze the selected articles based on performance metrics used for the evaluation of any data provenance technique?) and the application domain that the review article covers.
@@
Wang et al. (2016b) provide a deep understanding of provenance schemes, categorizing them into five main categories: distributed schemes, elementary schemes, lossy compression schemes, lossless compression schemes, and block schemes. The survey primarily focuses on WSN networks, examining the techniques used and discussing their respective advantages and disadvantages. One notable limitation of the study is the absence of considerations regarding attacks and performance metrics in relation to the reviewed works. Another survey that categorizes data provenance techniques is presented by Hu et al. (2020). The categorization is based on three main technologies logging-based, cryptography-based and blockchain-based technologies. The authors presented the provenance techniques based on the technology used and the requirements to achieve a secure provenance framework.  Zafar et al. (2017) conducted a comparative analysis in their work, presenting a detailed taxonomy of secure provenance schemes. They further perform a discussion of existing schemes, focusing on their strengths and weaknesses. The survey provides a deep understanding of the concept of provenance and its lifecycle, while addressing the necessary security requirements for such systems. The survey does not mention the performance metrics used in the studied methods.  Alam and Wang (2021) propose a comprehensive survey that present and analyze different research methods on three aspects that are provenance management, capture, and analysis. They additionally address the problems of maintaining data integrity, identifying attack chains in trigger-action platforms, and policy compliance in provenance systems of IoT environments.
@@
A comprehensive review on data provenance collection and security in distributed environments is presented by Ametepe et al. (2021). The work classifies provenance schemes based on provenance collection schemes, general provenance schemes, and basic provenance schemes. The paper discusses provenance security elements based on integrity, confidentiality and availability, but does not discuss provenance in IoT environemnt.  Gultekin and Aktas (2022) provide an SLR-based methodology on data provenance in IoT. The work selects 16 papers for the study. These papers are discussed based on the main subject that the study focus on. The survey lacks comparative analysis based on security requirements and other metrics that are needed to evaluate a provenance system. The most recent survey on data provenance is provided by  Pan et al. (2023). The study explores the significance of data provenance in the domains of security and privacy. Additionally, the authors outline the foundational principles and models of data provenance, while investigating the mechanisms proposed by previous studies to achieve security objectives. Furthermore, they conduct a comprehensive review of existing schemes aimed at securing the collection and manipulation of data provenance, commonly referred to as secure provenance and the role of data provenance, specifically in terms of threat provenance. The study covers various application domains beyond IoT, including file systems, databases, generic systems, cloud computing, data management, distributed networks, and media data. While the survey examines many of the requirements for provenance systems, it does not specifically address the performance evaluation, deployment technologies and storage methods of provenance techniques in IoT environments. Furthermore, it lacks a comparative analysis in this regard.
@@ 5. Security techniques for data provenance
The idea of data provenance has been applied by many researchers for identifying the source of data, trace ownership to ensure data authenticity, and evaluate trustworthiness. We classify and study the provided solutions based on the following main technologies: Cryptography-based, Bloom Filters, Link Fingerprints, Blockchain, Watermarking, Physical Unclonable Functions, graph-based provenance, data sanitization, lexical chaining, path difference, logging-based and frameworks that uses different storing methods. In Appendix A, we provide a detailed comparison of the selected papers based on a number of protocol metrics, performance metrics, security requirements, and attacks (cf. Tables 4, 6, 7, and 8, respectively).
5.1   Cryptography-based Techniques — Various cryptographic techniques, including symmetric and asymmetric cryptography methods, hash functions, and digital signatures, are applied to design and implement systems for identifying the source of data and ensuring the integrity of both data and provenance. These concepts are used to propose solutions for data provenance in many applications. The literature includes the largest number of these solutions when dealing with cryptographic approaches. The trustworthiness of the data items in a sensor network is evaluated by  Lim et al. (2010). They compute trust scores for sensor nodes and data packets by using data provenance. These scores provide a method of indicating the level of trustworthiness of both nodes and data items. This method provide security and trustworthiness for sensor networks, yet they fail to address the challenge of retrieving data provenance against various attacks. Moreover, a dictionary-based secure provenance approach for WSN is proposed by  Wang et al. (2016a). They embed path indexes in the provenance instead of the actual data path by using packet path dictionaries. Therefore, compared to the current lossy provenance systems, the compressed provenance size in the proposed lossless approach is smaller. To achieve security requirements such as integrity, availability and authenticity of provenance, the AM-FM sketch method and a robust packet sequence number generating technique are used in the system. Provenance records are encoded at nodes that engage in every stage of data processing and transmission using the suggested dictionary-based system. A secure message authentication code integrates the packet and its provenance together to provide security against any unauthorized change. After verifying the Message Authentication Code (MAC) during decoding, the BS retrieves the packet’s provenance graph.
While the majority of previous studies have concentrated on how to protect against data manipulation in Home Area Networks (HAN) networks,  Chia et al. (2017) introduce a security topic that has received less attention in such networks, namely data provenance. To ensure that the stated energy usage is actually consumed and is gathered from the specified node at the exact location, they provide a unique technique based on threshold cryptography and Shamir’s secret sharing (Shamir, 1979). The authors describe a unique use of secret sharing schemes for home energy monitoring networks, in which a secret, or a single private key, is distributed among all the members who provide data on energy consumption. This system also addresses the issue of location verification. To achieve this, authors incorporate a location generator as an additional component. The location generator employs trilateration techniques based on Received Signal Strength Indicator (RSSI) values and RSSI filtering methods to verify the location of the device. This not only ensures that the right power source is under observation but also detects any potential device relocation. Also,  Suhail et al. (2018) provide a solution to the challenge of integrating IoT with a system that is aware of data provenance, enabling the tracking of data flow across nodes and monitoring data transformations applied by nodes in the network. They introduce a lightweight method for transmitting provenance information for IoT sensor data. This method encodes data provenance information using a hash chain scheme as it traverses each participating node, with the final provenance verification taking place at the destination node. The technique is implemented within the context of the IPv6 Routing Protocol for Low-Power and Lossy Networks (RPL) (Alexander et al., 2012) in Contiki OS (Dunkels et al., 2004). Each generated data packet includes three fields: packet sequence number, data payload and provenance information (i.e. hash value of node ID). Furthermore,  Xu et al. (2019) present a provenance approach for WSN based on path index differences, where a new packet path is encoded using the index of a highly similar backbone path and the differences between them. Even if the topology of the WSN is unstable, this technique may reach an increased provenance compression rate compared to the dictionary based provenance scheme (Wang et al., 2016a). In the presented backbone path selection approach for WSN, the gradient of the node on the data packet path is also determined. In order to cover the commonly used packet transmission channels, the authors also develop a comparable backbone path elimination technique for WSN. To find the packet pathways with the highest similarity in the dictionaries, they develop a locality-sensitive hashing (LSH), SimHash (Buyrukbilen and Bakiras, 2014; Charikar, 2002), based similarity comparison algorithm. The proposed approach is evaluated according to Total Energy Consumption (TEC) and Average Provenance Size (APS).
In their work, Porkodi and Kesavaraja (Porkodi and Kesavaraja, 2021) present a framework to secure the data provenance by introducing blockchain and access control policies in IoT systems. They use hybrid attribute-based encryption to securely transmit provenance information. The proposed method is evaluated using different performance metrics such as computational cost and the throughput of encryption/decryption, and the key strength is calculated using the effect of avalanche. Additionally, the authors conducted experiments to prove that the proposed approach reduces computational cost and achieves high throughput. In their research,  Xu et al. (2022) introduce two provenance schemes for Wireless Sensor Networks. The first is the Path Index Differences-based Provenance (PIDP) scheme, where provenance information is encoded using differences in packet path indices. Specifically, it links an index representing the packet path’s similarity to the main backbone path and the variation between the actual packet path and this backbone path. Additionally, the authors propose a second scheme known as the Path Hash Value-based Provenance (PHP) scheme. In this approach, provenance data is encoded as a combination of the data source node’s ID and a part of the hash value taken from the packet’s path. Both schemes are evaluated based on their provenance compression rate and energy conservation rate, and these metrics are then compared to those of the DP scheme. Xu and Wang (Xu and Wang, 2022) introduce a provenance scheme called Multi granularity Graphs-based Stepwise Refinement Provenance (MSRP). In this scheme, they utilize mutual information between pairs of nodes as a similarity index to categorize node IDs. This categorization forms the basis for generating multi granularity topology graphs. Additionally, they apply the Dictionary-based Provenance (DP) scheme for stepwise encoding of the provenance information. The Base Station (BS) follows the same stepwise approach for provenance recovery and simultaneously conducts data trustworthiness evaluation during decoding. The performance of the MSRP scheme is thoroughly evaluated through a combination of simulations and testbed experiments. Results show that the scheme achieve high provenance compression rate, energy utilization and efficiency of data trustworthiness assessment.
5.2   Bloom Filters — A compact data structure, using hashing, enables rapid verification of item presence within the structure. Provenance information is embedded within the generated structure and the bloom filter is transmitted along with the data. By employing this method, the original information remains inaccessible to potential adversaries. Due to the variability of the bloom filter value from one packet to another, establishing a connection or association between previous bloom filters and recent data becomes challenging. In their work,  Sultana et al. (2015) present a secure provenance scheme for WSN. Their approach involves embedding provenance information into a Bloom filter (BF), which is transmitted alongside the data. This scheme effectively addresses the challenges posed by resource constraints in WSN. It requires a single channel for data and provenance transmission. The scheme overlooks data integrity and only focuses on studying the packet drop attack. Additionally,  Siddiqui et al. (2019) present a data provenance technique for IoT devices that employs Bloom Filter and attribute-based encryption. This approach presents challenges as IoT devices typically have limited memory capacity, making it impractical to store extensive provenance information. Furthermore, the technique is vulnerable to physical attacks, whereby an attacker can easily manipulate the stored provenance information within an IoT device.
Harshan et al. (2020) introduce a method for embedding provenance information known as the Deterministic Double-Edge (DDE) embedding mechanism, which is based on Bloom filters (Lu et al., 2012; Shebaro et al., 2012). In this approach, a relay node embeds information about both edges connected to it: the edge through which it receives a packet and the one through which it sends the packet. This means that a node can effectively cover two edges in the path with a single action, reducing the need for the next node in the path to update the provenance. This results in decreased packet delay. Authors show that the hop-counter within the provenance segment can also be utilized to coordinate the skipping strategy among nodes. Additionally, they propose upper bounds on the error rates of the DDE embedding technique, which are expressed as functions of the network’s node count, the number of hops, the size of the Bloom filter, and the number of hash functions computed by each node.
5.3   Fingerprints — Fingerprints refers to unique identifiers or signatures assigned to the communication links between IoT devices. These fingerprints are typically generated based on various characteristics or attributes of the link, such as received signal strength, latency, packet loss, or other network metrics. The homomorphic feature of public-key cryptography is exploited by numerous anonymous fingerprinting systems. These approaches enable the user’s fingerprint to be embedded in an encrypted domain (using public key) so that only the user may access the decrypted fingerprinted material by using the private key (Megías, 2015). Data provenance solutions have been developed by analyzing and comparing link fingerprints generated from different attributes. In their work,  Ali et al. (2014) present a method for enhancing the security of data provenance in bodyworn medical sensor devices. They achieve this by using the spatio-temporal characteristics of the wireless channels used for communication by these devices. This solution allows two parties to create highly similar link fingerprints, which uniquely link a data session to a specific wireless connection. This link enables a third party to later verify transaction details, especially the specific wireless link through which the data was transmitted. To validate this approach, experimental testing was conducted using MicaZ motes running TinyOS, and the results show an improvement in energy efficiency. Additionally, the proposed technique generates provenance information on each session, reducing the use of cryptographic methods. Using different fingerprint method, Alam and Fahmy (Alam and Fahmy, 2014) propose a provenance encoding and construction method that adapts three encoding schemes: juxtaposition of ranks, prime scheme, and Rabin fingerprints. The method is referred to as Probabilistic Packet Flow (PPF). In the first scheme, provenance is constructed based on the rank of the node instead of the node ID, which requires fewer bits for encoding. Assuming that the packet contains a specific space to hold the identities of $m$ nodes, a counter of $\log(m)$ bits is used to track the embedded ranks in the packet. The second scheme is based on prime multiplication, which embeds more node IDs using the same number of bits compared to the previous one. This method uses prime numbers as node IDs, and their multiplication results in encoding a set of IDs that can be uniquely factorized. In their third method, a data path traversed by a packet is a sequence of bits that represents the IDs of nodes on that path. The fingerprint of this sequence of bits is transmitted instead of the actual sequence.  Kamal and Tariq (2018) introduce a lightweight protocol for a multi-hop IoT network, aiming to ensure both data security and the establishment of data provenance. The protocol utilizes link fingerprints derived from the RSSI of IoT nodes within the network. The protocol achieves data provenance by appending the encoded link fingerprint to the data packet header as it traverses each node. After receiving the packet, the server decode the packet header in a sequential process. However, provenance data expands rapidly, which requires transmitting a large amount of provenance information to data packets, thereby increasing the bandwidth overheads.
5.4   Blockchain-based Solutions — Distributed ledger technologies, like blockchain, have gained visibility for maintaining the security and privacy of provenance data. These blockchain platforms serve as the primary means to ensure data integrity by creating transactions as provenance records that are chained in a secure manner. Once recorded on the blockchain, these records cannot be altered or removed, thereby establishing a trustworthy system for verifying the integrity of the provenance information in a network environment that is vulnerable to different attacks.
In their work,  Javaid et al. (2018) propose a solution for data integrity and data provenance in IoT networks by using physical unclonable functions (PUFs) and a blockchain variant with smart contracts that is Ethereum. This method is called BlockPro. PUFs are used to maintain data provenance by providing unique hardware fingerprints. To overcome data tampering attacks and block unregistered devices, Ethereum is used as a decentralized digital ledger. With the presence of expanding sequence of records, data undergoes initial validation before being permanently stored on the blockchain. Once stored, it can not be tampered with or modified, thus ensuring data integrity.  Sigwart et al. (2020) implement an IoT data provenance framework based on smart contracts using a generic data model to provide data provenance capturing, storing, and querying functionalities for different IoT use cases. This work extends their approach proposed in (Sigwart et al., 2019). The framework uses the data provenance model by  Olufowobi et al. (2017). The authors conduct an assessment of the proposed framework with specified requirements through the implementation of a proof-of-concept using Ethereum smart contracts. In this approach, a provenance record ($prov(dp)$) for a data point ($dp$) is composed of a 3-part structure. It links the address ($addr(dp)$) of $dp$ (i.e. essentially an identifier ID) with the set of provenance records associated with the data points that are used to create a $dp$ (referred to as $inputs(dp)$), and includes a context element ($context(dp)$). This element includes information for provenance purposes, such as details about the agents involved in the computation of the data point, timestamp, location, or the specific execution context within the IoT system.
@@
5.5   Watermarking and Physical Unclonable Functions — Watermarking is a widely known advancement in the field of WSN security. It serves the purpose of identifying any alterations made to sensory data, effectively preventing unauthorized interception (Megías et al., 2010; Araghi et al., 2023). The two main categories of digital watermarking are fragile watermarking and robust watermarking, each offering different anti-attack properties. Fragile watermarks become undetectable when data is modified, whereas robust watermarks can withstand various forms of distortion (Cox et al., 2007; Megías, 2021). Watermarking is being used in many security applications and is introduced in some proposed data provenance works.  Sultana et al. (2011) develop a technique for data provenance aimed at identifying malicious packet dropping attacks. The technique depends on the timing characteristics between packets after the process of embedding provenance information. Based on the distribution of inter-packet delays it detects the packet loss. Then, it identifies the presence of an attack and localizes the malicious node or link.
Similarly, Physical Unclonable Functions (PUFs) provide a hardware-oriented system that generates a response for a specific challenge, ensuring uniqueness for each device in the system. PUFs facilitate the authentication of IoT devices that generate sensitive data while maintaining device anonymity. Through the presence of a trusted third-party verifier, the identities of these devices can be verified without compromising their anonymity (Kanuparthi et al., 2013). Aman et al. (2019) developed two secure protocols for data provenance in IoT networks, aiming to achieve authentication and privacy preservation. These protocols address two different scenarios: the first scenario involves a direct connection between an IoT device and a wireless gateway, while the second scenario deals with IoT devices indirectly connected to the wireless gateway through multiple hops of other IoT devices. Both protocols use PUFs in addition to wireless link fingerprints generated from the RSSI between communicating nodes. Through experimental analysis, the authors show that these protocols achieve high efficiency in terms of computational complexity and energy utilization, while also obtaining robustness against various physical and cloning attacks. Aman et al. (2021) propose an analytical model to create a mechanism that enables the establishment of data provenance in IoT systems. Their approach incorporates PUFs and the extraction of fingerprints from the wireless channel, along with the implementation of mutual authentication and anonymity measures, all aimed at achieving robust data provenance. The approach lacks consideration for the multi-hop scenario and fails to adequately address tracking of data packet provenance. Moreover, Hamadeh and Tyagi (Hamadeh and Tyagi, 2021) propose an approach which combine two solutions: data provenance and privacy prserving in IoT networks. They provide trustworthiness and dependable IoT networks by using PUFs with non-interactive zero knowledge proof. In this method, an IoT device has the capability to transmit data to its respective server without revealing its identity, as it provides proof of ownership. In particular, the method under consideration is designed to validate that the authorized device executed an authorized program for creating or modifying data. Authors introduce a privacy-centric data provenance protocol. To validate practicality and effieciency, they developed the protocol using Altera Quartus and subsequently implemented it on an Altera Cyclone IV FPGA.
5.6   Graph-based Provenance — Graph-based provenance is a method used in data provenance encoding to represent the origin or history of data. These graphs capture the relationships and dependencies among data elements, processes, and transformations, providing a visual or structural representation of how data has evolved over time. Three different types of graph-based provenance representations as descibed below.
Event-flow graphs are a type of provenance graph that represents data provenance by capturing events and the flow of data between them. Each event corresponds to a data operation or transformation, and edges in the graph indicate the data flow from one event to another.  Chang et al. (2022) propose a data provenance approach for provenance systems in IoT applications based on event-action flows instead of provenance graphs. Even-action flow is defined as a sequence of events and actions including a time-stamp in an execution trace. They are less complex than provenance graphs due their simple structure form which allows users to understand it easier. In their work, the authors present an event-flow graph to regular users as a static abstraction of every possible provenance graph for IoT applications. They dynamically link time-stamped events and actions to statically create event and action nodes. Then, users can query provenance information from the event-flow graph by selecting an event or action node to choose the associated time-stamped actions or events. After users create a query by picking a timestamped event or action, the system will respond with one of two types of information: “What provenance” answers questions about which events or actions are triggered by the user’s specified action or event, and “Why provenance” provides insights into the events or actions responsible for causing the user-specified action or event. The system is developed as a Graphical User Interface (GUI) forming a user-friendly graphical provenance system.
Finally, the use of Directed Acyclic Graphs (DAG) allows directed structures with no cycles. This means that data transformations and dependencies are represented as a directed flow without loops or feedback. DAGs are used to represent provenance information because they are well-structured and provide efficient querying and tracking of provenance records. ProChain framework is a provenance-aware approach of traceability proposed by Al-Rakhami and Al-Mashari (Al-Rakhami and Al-Mashari, 2022) for IoT-based supply chain systems. The IOTA protocol, a third-generation DLT, is utilized by the ProChain framework. To overcome limitations and provide a scalable, quantum-resistant, and attack-proof solution for the systems built around IoT, it makes use of the DAG information structure in contrast to the linear structure used by the blockchain (Silvano and Marcelino, 2020). ProChain enables food item traceability from manufacturing to retailer with the use of several IoT sensors and provenance data at each engaged supply chain phase. ProChain strengthens and improves the management and optimization of all operations while also serving as a guarantee for the quality and safety of food. On the Raspberry Pi 3B platform, the ProChain idea is evaluated by simulating an IoT-deployed supply chain. The average measured time and energy consumption were then evaluated to check on the usability of the framework. The authors review and implement the framework to show how it may be used in supply chain systems to couple supply chain data with the IOTA Tangle and generate provenance information by adding data for various payload sizes.
@@
5.7   Sanitization and Lexical Chaining — There are situations where privacy concerns prevent some information from being shared inside the provenance chain. This is where the method of data sanitization is presented. Data sanitization assists in hiding some important information which could compromise privacy. Data sanitization is mainly used to make sure that data is consistent, accurate, and dependable, such that it can be used for reporting, analysis, and data-hiding for privacy preservation.  Lomotey et al. (2019) present two approaches for device and data verification in an IoT network to achieve trust and privacy preservation. The first approach enables devices to subscribe and reveal their metadata to allow for data packet tracing without compromising privacy. To achieve this, a data sanitization method for hiding sensitive information is applied. Users and devices have the option to label attributes within the provenance data as non-shareable with other devices in the chain. In the second approach, the authors modeled the entire peer-to-peer IoT network as a graph network. Data origins are verified using Floyd’s algorithm between different interconnected nodes. This work ensures transparency, traceability, and privacy preservation through both proposed approaches.
@@
Semantically similar words or phrases are linked together by lexical cohesiveness. The related elements can be linked together to create Lexical Chains after all cohesive relationships have been determined forming a conceptually accurate building blocks in a variety of natural language processing systems (Remus and Biemann, 2013). This is how a typical chain may be written:
@@
would be feasible to track and categorize the communication channels between different IoT devices as well as their previous interactions using this method, which may be used for machine-to-machine communication. In their work,  Lomotey et al. (2018) first emphasize the use of provenance through the design of algorithms to confirm the origin of IoT-based data, and then they propose developing completeness techniques through visual analytics to trace data packets through the complete route in an IoT network. They present an improved system provenance mechanism in their study to achieve traceability. The technique is based on an associative logic to decide all connections established in a machine-to-machine communication between IoT nodes. To create links between device communication and data propagation to the $n$-th degree, a statistical lexical chaining based on the Adjusted Rand Index (ARI) is suggested as an alternative to knowledge-based methodology. Because data propagation pathways and object-to-object communications can be identified, the proposed IoT architecture makes traceability easier. Based on their findings, the suggested system demonstrates that, in terms of identifying linkability, unlinkability, and availability, the ARI is more accurate than the knowledge-based methodology. Additionally, visual analytics is offered to give a clearer understanding of interconnection in IoT nodes using visualization graphs such as HyperTree Graph, the Weighted Graph, and the combination of SpaceTree and RGraphs.

### v2

@@ 5. Security Attacks
IoT networks are vulnerable to active and passive attacks. The proposed security techniques takes into account these attacks and try to prove their robustness against it through many proposed solutions using different technologies. Provenance information may be more sensitive than the data itself, which makes it a priority for attackers. Attacks on such systems are of two types: attacks on data and attacks on provenance. Next, we outline some representative attacks that may be perpetrated in an IoT environment against data and provenance.
5.1. Data Attacks
@@
  • Packet drop attack: It is a security threat where an adversary intentionally discards or drops specific data packets within the network. This malicious act disrupts communication between IoT devices and can lead to various issues, such as data loss, delayed information delivery, or service unavailability. As IoT devices often rely on wireless communication and may operate in resource-constrained environments, packet drop attacks can significantly impact the overall system’s functionality and reliability.
@@
  • Packet replay attack: It is a type of security attack where an adversary intercepts and records data packets as they travel through the network. The attacker then replays or re-sends these recorded packets at a later time, attempting to deceive the system into thinking the information is fresh and legitimate. This malicious action can lead to various security issues, such as unauthorized access, data manipulation, or false triggering of actions based on the replayed data.
@@
  • Data forgery: It is a form of security attack where an adversary manipulates or alters the data transmitted between IoT devices. The attacker may forge false sensor readings, control commands, or other critical information to deceive the system or cause malicious actions. This type of attack can lead to incorrect decisions, unauthorized access, or compromised system integrity.
@@
  • Data modification attack: It is a type of security attack where an unauthorized entity alters or modifies the data being transmitted between IoT devices. The attacker may tamper with sensor readings, control commands, or other critical information, leading to incorrect decisions or actions within the system. This can result in operational disruptions, potential risks, or unauthorized access to sensitive data.
@@
  • Eavesdrop: It refers to the unauthorized interception and monitoring of communication between IoT devices. A malicious attacker, also known as an eavesdropper, captures and listens to data packets as they traverse the network. By doing so, the eavesdropper can access sensitive information, such as sensor readings, control commands, or personal data, which can lead to privacy breaches, security vulnerabilities, and potential misuse of the intercepted data.
5.2. Provenance Attacks
@@
The studied security techniques in the literature, along with the types of attacks each method addresses, are summarized in Table 16 of Appendix A. We also provide a taxonomy for these attacks as shown in Figure 3. According to Table 16, it is evident that the majority of the reviewed papers do not consider most types of attacks. While most papers address forgery and modification attacks, both for data and provenance, they tend to overlook other significant attack vectors such as replay attacks, packet drop, and provenance chain tampering. This limited focus on specific types of attacks suggests that the current security solutions are not fully robust, leaving room for future research to explore more holistic approaches that account for a wider range of potential threats in IoT networks.
@@ 6. Existing surveys
Several review articles address the issue of data provenance in IoT networks. Most of these articles focus on the security requirements for data provenance in IoT. In this section, we compare and highlight the importance and need to conduct this review paper. Table 4 provides a comparison between the different review articles based on a number of aspects, including time range (represent the starting and ending years of the conduct review), review methodology (Do the article follow a specified review methodology to conduct the review?), taxonomy (Does the authors provide researchers with a taxonomy that summaries the different aspects used in the review?), security requirements (Does the review use and elaborate the needed security requirements in data provenance approaches?), attacks (Does the review analyze the different approaches based on the security attacks?), performance metrics (Does the review analyze the selected articles based on performance metrics used for the evaluation of any data provenance technique?) and the application domain that the review article covers.
@@
Wang et al. (2016b) provide a deep understanding of provenance schemes, categorizing them into five main categories: distributed schemes, elementary schemes, lossy compression schemes, lossless compression schemes, and block schemes. The survey primarily focuses on WSN networks, examining the techniques used and discussing their respective advantages and disadvantages. One notable limitation of the study is the absence of considerations regarding attacks and performance metrics in relation to the reviewed works. Another survey that categorizes data provenance techniques is presented by Hu et al. (2020). The categorization is based on three main technologies logging-based, cryptography-based and blockchain-based technologies. The authors presented the provenance techniques based on the technology used and the requirements to achieve a secure provenance framework.  Zafar et al. (2017) conducted a comparative analysis in their work, presenting a detailed taxonomy of secure provenance schemes. They further perform a discussion of existing schemes, focusing on their strengths and weaknesses. The survey provides a deep understanding of the concept of provenance and its lifecycle, while addressing the necessary security requirements for such systems. The survey does not mention the performance metrics used in the studied methods.  Alam and Wang (2021) propose a survey that present and analyze different research methods on three aspects that are provenance management, capture, and analysis. They additionally address the problems of maintaining data integrity, identifying attack chains in trigger-action platforms, and policy compliance in provenance systems of IoT environments.
@@
A review on data provenance collection and security in distributed environments is presented by Ametepe et al. (2021). The work classifies provenance schemes based on provenance collection schemes, general provenance schemes, and basic provenance schemes. The paper discusses provenance security elements based on integrity, confidentiality and availability, but does not discuss provenance in IoT environemnt.  Gultekin and Aktas (2022) provide an SLR-based methodology on data provenance in IoT. The work selects 16 papers for the study. These papers are discussed based on the main subject that the study focus on. The survey lacks comparative analysis based on security requirements and other metrics that are needed to evaluate a provenance system. The most recent survey on data provenance is provided by  Pan et al. (2023). The study explores the significance of data provenance in the domains of security and privacy. Additionally, the authors outline the foundational principles and models of data provenance, while investigating the mechanisms proposed by previous studies to achieve security objectives. Furthermore, they conduct a review of existing schemes aimed at securing the collection and manipulation of data provenance, commonly referred to as secure provenance and the role of data provenance, specifically in terms of threat provenance. The study covers various application domains beyond IoT, including file systems, databases, generic systems, cloud computing, data management, distributed networks, and media data. While the survey examines many of the requirements for provenance systems, it does not specifically address the performance evaluation, deployment technologies and storage methods of provenance techniques in IoT environments. Furthermore, it lacks a comparative analysis in this regard.
@@ 7. Security techniques for data provenance
The idea of data provenance has been applied by many researchers for identifying the source of data, trace ownership to ensure data authenticity, and evaluate trustworthiness. We classify and study the provided solutions based on the following main technologies: Watermarking, Data Sanitization, Lexical Chaining, Path Difference, Logging-based Techniques, Bloom Filters, Fingerprints, Frameworks using Storing Methods, Graph-based Provenance, Cryptography-based Techniques, Physical Unclonable Functions, and Blockchain-based Solutions. In this section, the provenance encoding techniques are ordered by increasing overhead, beginning with lighter methods and progressing to those with higher resource demands. In Appendix A, we provide a detailed comparison of the selected papers based on a number of protocol metrics, performance metrics, security requirements, and attacks.
@@
5.1   Watermarking — Watermarking is a widely known advancement in the field of WSN security. It serves the purpose of identifying any alterations made to sensory data, effectively preventing unauthorized interception (Megías et al., 2010; Araghi et al., 2023). The two main categories of digital watermarking are fragile watermarking and robust watermarking, each offering different anti-attack properties. Fragile watermarks become undetectable when data is modified, whereas robust watermarks can withstand various forms of distortion (Cox et al., 2007; Megías, 2021). Watermarking is being used in many security applications and is introduced in some proposed data provenance works. As shown in Table 5, one work discusses watermarking as a solution for data provenance protection.  Sultana et al. (2011) develop a technique for data provenance aimed at identifying malicious packet dropping attacks. The technique depends on the timing characteristics between packets after the process of embedding provenance information. Based on the distribution of inter-packet delays it detects the packet loss. Then, it identifies the presence of an attack and localizes the malicious node or link.
@@
5.2   Data Sanitization — There are situations where privacy concerns prevent some information from being shared inside the provenance chain. This is where the method of data sanitization is presented. Data sanitization assists in hiding some important information which could compromise privacy. Data sanitization is mainly used to make sure that data is consistent, accurate, and dependable, such that it can be used for reporting, analysis, and data-hiding for privacy preservation.  Lomotey et al. (2019) present two approaches for device and data verification in an IoT network to achieve trust and privacy preservation. The first approach enables devices to subscribe and reveal their metadata to allow for data packet tracing without compromising privacy. To achieve this, a data sanitization method for hiding sensitive information is applied. Users and devices have the option to label attributes within the provenance data as non-shareable with other devices in the chain. In the second approach, the authors modeled the entire peer-to-peer IoT network as a graph network. Data origins are verified using Floyd’s algorithm between different interconnected nodes. This work ensures transparency, traceability, and privacy preservation through both proposed approaches. The technique is presented in Table 5.
@@
5.3   Lexical Chaining — Semantically similar words or phrases are linked together by lexical cohesiveness. The related elements can be linked together to create Lexical Chains after all cohesive relationships have been determined forming a conceptually accurate building blocks in a variety of natural language processing systems (Remus and Biemann, 2013). This is how a typical chain may be written:
@@
would be feasible to track and categorize the communication channels between different IoT devices as well as their previous interactions using this method, which may be used for machine-to-machine communication. In their work,  Lomotey et al. (2018) first emphasize the use of provenance through the design of algorithms to confirm the origin of IoT-based data, and then they propose developing completeness techniques through visual analytics to trace data packets through the complete route in an IoT network. They present an improved system provenance mechanism in their study to achieve traceability. The technique is based on an associative logic to decide all connections established in a machine-to-machine communication between IoT nodes. To create links between device communication and data propagation to the $n$-th degree, a statistical lexical chaining based on the Adjusted Rand Index (ARI) is suggested as an alternative to knowledge-based methodology. Because data propagation pathways and object-to-object communications can be identified, the proposed IoT architecture makes traceability easier. Based on their findings, the suggested system demonstrates that, in terms of identifying linkability, unlinkability, and availability, the ARI is more accurate than the knowledge-based methodology. Additionally, visual analytics is offered to give a clearer understanding of interconnection in IoT nodes using visualization graphs such as HyperTree Graph, the Weighted Graph, and the combination of SpaceTree and RGraphs. These methods are presented in Table 5.